OpenStudy (anonymous):

Differential Equations: For what domain does y = x^2 satisfy the differential equation xy' = 2y Book says answer, for all x except x = 0, but why is this if y is define for all x including 0??? book error?

OpenStudy (anonymous):

When you separate the variables, you put x and y into the denominator; clearly they can't be zero in that position.

OpenStudy (anonymous):

i dont understand what you mean by seperate the variables... is that some kind of method? because if so I didnt get to that yet so I dont see why the book would ask a question on something is didnt cover yet o_O what i got is \[y=x^2\] \[y' = 2x\] substituing in the DE: \[x(2x) = 2(x^2)\] \[2x^2 = 2x^2\]

OpenStudy (amistre64):

it being a solution is not the issue; the domain for the solution is the issue

OpenStudy (anonymous):

but y and y' are define for all x so whats the problem, how does x get in the denominator?

OpenStudy (anonymous):

sorry i literally just started DE's today

OpenStudy (amistre64):

then itll make more sense to you later :) how do we find y=2x to be a solution?

OpenStudy (amistre64):

and by that i meant y = x^2 :)

OpenStudy (amistre64):

xy' = 2y y' - 2y = 1/x e^-2x y' - 2e^-2x y = e^-2x 1/x integrate: e^-2x y' - 2e^-2x y = e^-2x 1/x e^-2x y = integrate (e^-2x)/x dx y = integrate (e^-2x)/x dx ------------------- e^-2x

OpenStudy (amistre64):

at any rate; you see that we get an x in a denominator position

OpenStudy (anonymous):

not sure why you integrating the DE but I guess you're right and it will make sense later. So far the only definition the book gave for it being a solutions is this: http://i.imgur.com/dGxWh.png and there is no mention of integration but I guess ill move on and come back to this later when I get better understanding

OpenStudy (anonymous):

An alternate method.....\[xy' = 2y \implies x \frac{dy}{dx}=2y \implies \frac{dy}{y}=2\frac{dx}{x}\implies \ln(y)=2\ln(x)+C=\ln(x^2)+C\]\[\implies y=C_1x^2\]Clearly if the constant C1 is one, we get the desired solution.